# Predictive Linear Online Tracking for Unknown Targets

### v1

@@ 5.1.1 Target with a $V_{T}=0$ path length
Given the dynamic regret analysis and the main result in Theorem 4.1, the smaller the path length $V_{T}$ the easier the control task is for PLOT. To show its tracking performance, as well as the effect of the prediction horizon on such a target, we consider the circular target detailed in Example B.2 that has a $V_{T}=0$ path length. With this target revealed and measured online, as described in Section 2, we run PLOT repeatedly for various horizon lengths. The trajectory plots of the online target and the quadrotor are shown in Figure 1 for the first $T=2,3,5$ and $T=7$ seconds of the control. The regret plots for the considered multiple horizon lengths are shown in Figure 2 for a simulation of $T=200$ seconds.
@@
Firstly, both figures show that for too small $W$-s, PLOT exhibits poor tracking and regret performance. As expected from Corollary 4.2, PLOT with longer prediction horizons has better tracking by learning the reference dynamics online, achieving sublinear regret as verified by Figure 2. Additionally, the regret is increased for larger $W$-s up to a certain level. This behavior is explained by Theorem 4.1. The second term in the upper bound increases for larger $W$, however, is counteracted by the exponential constant $\rho^{W}$. See Appendix E.2 for further details.
@@ 5.1.2 Comparison with Other Online Control Methods
Next, we compare the dynamic regret of PLOT to other controllers. Namely, the disturbance-action policy (DAP) with memory proposed in Agarwal et al. (2019a), the Riccatitron algorithm from Foster and Simchowitz (2020), the SS-OGD algorithm by Karapetyan et al. (2023), the Follow the Leader (FTL) algorithm Abbasi-Yadkori et al. (2014), as well as the Naive LQR controller that does optimal state error feedback without any affine term. The reference target is chosen to follow the dynamics
@@
Figure 3 shows the accumulated regret of all algorithms. For the given challenging example, PLOT outperforms the others. This is mainly because, unlike all the others, PLOT implements a dynamic approach with a forgetting factor adapting to the fast-changing reference on time. Moreover, it deploys an indirect approach of learning the dynamics of the reference and then using it in the control, as opposed to a direct approach of learning the affine control term, implemented by all the others apart from LQR. For the given dynamic target the direct and static approaches produce a much smaller affine term compared to PLOT, leading to a performance very close to that of Naive LQR. A more detailed analysis of the comparison and implementation details is provided in Appendix E.3.
@@ 5.2 Experimental Validation on Quadrotors
Next, we validate the proposed algorithm on an actual quadrotor. Due to the non-linear dynamics of the quadrotor and the noisy state measurements, the hardware experiments have tight requirements for the stability and robustness of the control algorithm. In the hardware experiments, we define the states and actions the same as in the simulations. A laptop receives the state measurements, runs the online controller, and sends the control commands to the quadrotor via a radio communicator. The virtual reference trajectory is generated online after the drone has successfully taken off and is at a predefined hovering position. The controller receives the target at a rate of $10$ Hz during flight.
We implement the PLOT algorithm for the linearized model of the drones derived in Appendix E with a fixed prediction horizon of $W=5$ and a forgetting factor of $\gamma=0.8$. Figure 4 shows the trajectory plots of PLOT and naive LQR for a horizon of $T=40$ seconds. The naive LQR controller exhibits a delayed tracking behavior as expected, while PLOT achieves a smaller tracking error. Further details on the practical implementation of the algorithm are provided in Appendix F.
Figure 4: Trajectory plot of an ”Infinity”-shape reference tracked by the drone with the PLOT and LQR controllers.
@@ Acknowledgement
Appendix A Linear Quadratic Control Properties
@@ A.1 LQT properties
Let us now introduce the following proposition, which shows that the closed loop matrix $A-BK$ under the feedback gain $K$ defined in (8) is stable, that is, all of its eigenvalues are inside the unit circle. It also shows that the feedforward gains $K_{t},\forall t\leq T$ defined in (9) decay exponentially fast as $t$ increases. Under the optimal control law (7), we do not have stability in the classical sense; since the target state can be arbitrary the tracking error might not remain close to the origin. However, we have internal stability since all signals remain bounded as long as $r_{t}$ is bounded. A more general case with time-variant costs of the proposition is proven in Corollary 1 of Zhang et al. [2021b].
Proposition A.1 (Stability Zhang et al. [2021b]).
@@
$\|(A-BK)^{t}\|\leq\sqrt{\frac{\lambda_{max}(P)}{\lambda_{min}(P)}}\rho^{t},\,\rho=\sqrt{1-\frac{\lambda_{min}(Q)}{\lambda_{max}(P)}}\in(0,1)$
where $\lambda_{min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the minimum and maximum eigenvalue respectively, $P$ is defined in (6). As a result, the coefficient matrices $K_{t}$ defined in (9) satisfy
@@
for $c_{0}=\frac{\lVert B\rVert\lambda_{max}(P)}{\lambda_{min}(R)}\sqrt{\frac{\lambda_{max}(P)}{\lambda_{min}(P)}}$.
@@
Note that a pair $(Q,A)$ is detectable if and only if $(A^{\top},Q^{\top})$ is stabilizable. Under the above assumption, we still retain stability of $(A-BK)$ Anderson and Moore [2005]. The only difference is that we do not have an explicit characterization of $\rho$ anymore as in Proposition A.1.
A.2 LQT as LQR with disturbances
We can recast the LQT problem into an LQR problem with adversarial disturbances, see, for example, Karapetyan et al. [2023]. By redefining the tracking error as the system state $e_{t}=x_{t}-r_{t}$ and encoding the time-varying target states as the disturbance $w_{t}=Ar_{t}-r_{t+1}$, the resulting LQR problem is given as,
$\min_{u_{0:T-1}}$ $\sum_{t=0}^{T-1}\left(\lVert e_{t}\rVert_{Q}^{2}+\lVert u_{t}\rVert_{R}^{2}\right)+\lVert e_{T}\rVert_{P}^{2}$ (15)
@@ A.2 LQT as LQR with disturbances
Appendix B Autoregressive Systems
@@
B.1 Representation and learning complexity
@@
Example B.1 (Constant Velocity Target).
@@
Example B.2 (Circular Target with Constant Speed).
@@
In many previous works Karapetyan et al. [2023], Li et al. [2019], Nonhoff et al. [2023], learning complexity is captured by the total variation (path length) of the target itself
@@
In the case of ARX models we just replace $S_{t}$ with $\tilde{S}_{t}$. As shown in [Li et al., 2019, Th. 3], in the case of unstructured targets, e.g. randomly generated $r_{t}$, the former notion of complexity is optimal. However, we argue that in the case of dynamic structure, the former notion of complexity might be suboptimal. Consider the circular target or the linear target example. The total variation of the target state is linear with $T$ since $r_{t}$ is constantly changing: $L_{T}=\Omega(T)$. However, the total variation of the target dynamics is zero $V_{T}=0$. Using the PLOT algorithm would give us logarithmic regret in this case, while previous methods would give us linear regret.
We stress that every target can be represented trivially by ARX models. We can just set $S_{t}=0$, $v_{t}=r_{t}$. In this case, the total variation of $\tilde{S}_{t}=\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ is equal to $L_{T}$, as in prior work. However, the representation is non-unique in general. In many cases, the underlying dynamic structure will imply that a lower complexity representation exists, e.g., see Examples B.1B.2. Hence, our dynamic regret bounds supersede the ones in prior work.
B.2 Multi-step ahead dynamics expressions

### v3

@@ 5.1.1 Target with a $V_{T}=0$ path length
Given the dynamic regret analysis and the main result in Theorem 4.1, the smaller the path length $V_{T}$ the easier the control task is for PLOT. To show its tracking performance for such a target, as well as to visualize the effect of the prediction horizon $W$ on the control performance, we consider the circular target detailed in Example C.2 that has $V_{T}=0$ path length. With this target revealed and measured online, as described in Section 2, we run PLOT repeatedly for various horizon lengths. The trajectory plots of the online target and the quadrotor are shown in Figure 1 for the first $T=2,3,5$ and $T=7$ seconds. The regret plots for the considered multiple horizon lengths are shown in Figure 2 for a simulation of $T=200$ seconds.
@@
Firstly, both figures show that for too small $W$-s, PLOT exhibits a poorer tracking and regret performance due to the horizon truncation error. As expected from Corollary 4.2, given the nature of the reference target, PLOT with longer prediction horizons performs better by learning the reference dynamics online, achieving sublinear regret as verified by Figure 2. Additionally, the regret is increased for larger $W$-s only up to a certain saturated level. This behavior, as discussed in Section 4, is thanks to the multi-predictor setup of PLOT, making its regret scale with $\tilde{W}$, as opposed to linearly with $W$. See Appendix F.2.1 for further details.
5.1.2 Target with a $V_{T}=\mathcal{O}(\sqrt{T})$ Path Length
For more challenging targets, one can make use of Corollary 4.2 to tune for the best forgetting factor $\gamma$ if the path length $V_{T}$ is known. Here, we consider the case when only the order of the path length of a given reference is known, and, using Corollary 4.2, fix $\gamma_{a}=1-c_{\gamma}T^{-a}$, where $c_{\gamma}\in\mathbb{R}_{+}$ is a constant independent of $V_{T}$ and $T$, while $a\in\mathbb{R}_{+}$ is such that $\gamma_{a}\in(0,1]$ and is tuned based on $V_{T}$ and $T$.
We fix a spiraling reference target with $V_{T}=\mathcal{O}(\sqrt{T})$, with exact dynamics detailed in Appendix F.2.2. This order of $V_{T}$ results in the optimal value of $a=0.25$ and a forgetting factor $\gamma_{0.25}=0.78$ from Corollary 4.2, with $c_{\gamma}=1.50$ obtained experimentally.
PLOT’s performance with three difference forgetting factors is shown in the trajectory plot in Figure 3 for $T=40$ seconds, showing a better tracking performance for the regret-optimal $\gamma_{0.25}=0.78$. To show that $c_{\gamma}$ is independent of $T$, we fix, $c_{\gamma}$ to the tuned value of $1.50$ and perform experiments with varying horizon lengths from $T=150$ to $T=300$ seconds. For each $T$, PLOT is run with $4$ different forgetting factors with $a=0.1,~{}0.25,~{}0.5$ and $1$ (corresponding to $\gamma=1$), and the regret at the end of each experiment is shown in Figure 4. The results confirm that $\gamma_{0.25}=1-1.5T^{-0.25}$ consistently outperforms other forgetting factors in terms of regret, as expected from Corollary 4.2. Further details on this example are provided in Appendix F.2.2.
Figure 4: Regret of PLOT with varying $\gamma_{a}=1-c_{\gamma}T^{-a}$; $\gamma_{0.25}$ is regret-optimal based on Corollary 4.2.
@@ 5.1.3 Comparison with Other Online Control Methods
Next, we compare the dynamic regret of PLOT to other controllers. Namely, the disturbance-action policy (DAP) with memory proposed in (Agarwal et al., 2019a), the Riccatitron algorithm from (Foster & Simchowitz, 2020), the SS-OGD algorithm by (Karapetyan et al., 2023), the Follow the Leader (FTL) algorithm (Abbasi-Yadkori et al., 2014), as well as the Naive LQR controller that applies only an optimal state error feedback input without any affine term. The reference target is chosen to follow the dynamics
@@
where $s_{0}=1,\theta_{0}=0.06$, $T_{s}=0.1s$ and $s_{k}=-s_{k-1}$, $\theta_{k}=-0.99\times\theta_{k-1}$ for every $k=\sqrt{T}$ given some $T$ that we simulate $T=200$ seconds.
@@
Figure 5 shows the accumulated regret of all algorithms. For the given challenging example, PLOT outperforms the others. This is mainly because, unlike all the others, PLOT implements a dynamic approach with a forgetting factor adapting to the fast-changing reference on time. Moreover, it deploys an indirect approach of learning the dynamics of the reference and then using it in the control, as opposed to a direct approach of learning the affine control term, implemented by all the others apart from Naive LQR. For the given dynamic target the direct and static approaches produce a much smaller affine term compared to PLOT, leading to a performance very close to that of Naive LQR. The indirect approach of PLOT, however, may perform worse in cases where the reference targets lack dynamics and are mutually independent at each time step, as discussed in Remark 4.3. Further details on this, as well as the considered algorithms’ implementation are provided in Appendix F.3.
@@ 5.2 Experimental Validation on Quadrotors
We validate the proposed algorithm on the Crazyflie 2.1 quadrotor. Due to the non-linear dynamics of the quadrotor and the noisy state measurements, the hardware experiments have tight requirements for the stability and robustness of the control algorithm. In the hardware experiments, we define the states and actions to be the same as in the simulations. The control inputs are sent to the quadrotor through a radio communicator from a centralized computer that receives state measurements through a local area network connection to the motion capture system, and runs the online controller. The virtual reference trajectory is generated online after the drone has successfully taken off and is at a predefined hovering position. The controller receives the target at a rate of $10$ Hz during flight.
We implement PLOT for the linearized model of the drones derived in Appendix F with a fixed prediction horizon of $W=5$ and a forgetting factor of $\gamma=0.8$. Figure 6 shows the trajectory plots of PLOT and Naive LQR for a horizon of $T=40$ seconds and an “infinity”-shaped reference target.. The Naive LQR controller exhibits a delayed tracking behavior as expected, while PLOT achieves a smaller tracking error. Further details on the practical implementation of the algorithm are provided in Appendix G.
Figure 6: Trajectory plot of an “infinity”-shaped reference tracked by the drone with the PLOT and Naive LQR controllers.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Additional Related Work
Online and non-stochastic control.
Online and non-stochastic control techniques have been applied to systems with known (Agarwal et al., 2019a) as well as unknown system dynamics (Simchowitz, 2020). In this paper, we focus on the case of known system dynamics, where the main purpose of the online controller is to react to (non-stochastic) unknown disturbances/targets. [^2] Note that as shown in many works, see (Foster & Simchowitz, 2020; Karapetyan et al., 2023), the problem of tracking is equivalent under certain conditions to the problem of disturbance rejection. Hence, for the purpose of this literature review, we can use disturbance/target interchangeably. In this setting, Agarwal et al. (2019a) achieved $\mathcal{O}(\sqrt{T})$ regret for general convex stage cost functions under a disturbance affine feedback policy. In the case of quadratic costs, and more generally strongly convex stage costs [^3] Certain additional conditions are required., the guarantees were improved to logarithmic (Foster & Simchowitz, 2020; Simchowitz, 2020; Agarwal et al., 2019b). The aforementioned works consider static problems in the sense that they compete against the best stationary policy. The dynamic setting was studied in (Zhao et al., 2022), where dynamic regret guarantees of the order of $\mathcal{O}(\sqrt{TV_{T}})$ for general convex costs were derived for the first time, using gradient-based (first order) base learners. In (Baby & Wang, 2022) dynamic regret guarantees for the quadratic cost case (LQR) were derived using a variant of the Follow the Leading History (FLH) algorithm. It was proved that the dynamic regret with respect to any dynamic disturbance affine policy is upper bounded by $T^{1/3}V^{2/3}_{T}$, where $V_{T}$ is the total variation of the parameters of the policy. Learning disturbance affine feedback policies follows the direct online control paradigm since we adapt the policy parameters directly.
In this paper, we study implicit online control architectures, where we decouple online prediction of disturbances/targets from control design. Several works have studied the effect of predictions on receding horizon control (Yu et al., 2020; Li et al., 2019; Zhang et al., 2021b). However, either no method of prediction is provided or the dynamic regret guarantees are suboptimal in our setting. In particular, the regret guarantees scale with the total variation of the disturbance/target itself $\sum_{t=0}^{T-1}\|r_{t+1}-r_{t}\|$ (Li et al., 2019), which is appropriate for unstructured or slowly-varying targets but is suboptimal in the case of targets with dynamic structure; there exist targets for which the total variation of their state $r_{t}$ is linear, while the total variation of the target dynamics $S_{t}$ is zero–see Appendix C.1.
Abbasi-Yadkori et al. (2014) introduced the problem of non-stochastic control for target tracking, where the target can be adversarial. Logarithmic regret guarantees are provided for learning affine control policies with constant affine terms. However, policies with constant affine terms can only track static points as targets; they are not rich enough to track trajectories that are not points and, in general, time-varying targets. Recently, Niknejad et al. (2023) studied online learning for target tracking in the case of unknown system dynamics and unknown time-invariant target dynamics. Our setting is different since we have known dynamics for the system, time-varying dynamics for the target, and we use dynamic regret. Their static regret guarantee of $T^{2/3}$ is conservative for our setting, where we can obtain logarithmic regret in the static case, however, it applies to general convex costs and unknown dynamics. In (Karapetyan et al., 2023), dynamic regret guarantees are provided for target tracking, under a first order method which treats the targets as slowly varying or steady-state. Similar to (Li et al., 2019), the dynamic regret depends on the total variation of the target state which can be suboptimal in the case of target with dynamics. Zhang et al. (2022) proved $\tilde{\mathcal{O}}(\sqrt{|\mathcal{I}|})$ (for any interval $\mathcal{I}$) adaptive regret guarantees for tracking in the case of convex costs. Their setting is different since no dynamic structure is assumed for the targets, the comparator class is controllers which are constant over the interval of interest, while we focus on square losses and dynamic regret.
Although related, our setting is different from (Gradu et al., 2023) and (Minasyan et al., 2021) where the system itself is linear time varying, there is no target, and adaptive regret is considered as a performance metric. Instead, here, the system’s dynamics are time invariant, the target’s dynamics are unknown and time varying, and we consider dynamic regret. Safe online non-stochastic control subject to constraints has also been studied before (Zhou et al., 2023; Nonhoff et al., 2023; Li et al., 2021). Dealing with constraints is a more challenging problem in general. In many cases, the analysis is restricted to the static case (Li et al., 2021), requires relaxing the notion of policy regret (Zhou et al., 2023) or the path length scales with the target state (Nonhoff et al., 2023), which is not always optimal in our setting as we discuss above.
Finally, we note that when the system dynamics are unknown, it is no longer possible to achieve logarithmic regret in the general case, even under stochastic disturbances (Simchowitz & Foster, 2020; Ziemann & Sandberg, 2022). It is possible to achieve $\tilde{\mathcal{O}}(\sqrt{T})$ regret (Chen & Hazan, 2021; Simchowitz, 2020; Mania et al., 2019) instead.
Prediction and online least squares.
Online linear regression and the least squares method, in particular, have been studied extensively before (Azoury & Warmuth, 2001; Cesa-Bianchi & Lugosi, 2006). Anava et al. (2013) provided logarithmic regret guarantees for online prediction of auto-regressive (AR) and auto-regressive with moving average (ARMA) systems using the online Newton step (ONS) method. The result holds for exp-concave prediction loss functions, which includes the square prediction loss. Similar guarantees were proved for partially observed stochastic linear systems in the case of Kalman filtering (Ghai et al., 2020; Tsiamis & Pappas, 2022). Under first-order methods, the regret guarantees become $\mathcal{O}(\sqrt{T})$ (Anava et al., 2013; Kozdoba et al., 2019), however, first-order methods are also known to adapt to changes (Zinkevich, 2003; Hazan & Seshadhri, 2009). On the contrary, the standard versions of second-order methods like RLS or ONS are not adequate for time-varying settings, where adaptation is needed. Another challenge is that in the case of multi-step ahead prediction, these methods should be adapted to deal with delayed learning feedback.
To solve the former issue, one way is to employ multiple learners along with a meta-algorithm that chooses the best expert online, also known as the Follow the Leading History (FLH) algorithm (Hazan & Seshadhri, 2009; Baby & Wang, 2022). Another option is to employ exponential forgetting (Yuan & Lamperski, 2020; Ding et al., 2021). In this paper, we follow the second approach since it has been standard practice in the adaptive control literature (Guo & Ljung, 1995; Åström et al., 1977). Moreover, it has a simple implementation for real-time applications, like drone control.
Control Literature.
Common feedback control policies, such as the proportional-integral-derivative (PID) controllers find extensive use for tracking problems in practice (Pan et al., 2018; Cervantes & Alvarez-Ramirez, 2001). For example, integral control has been used to track constant unknown offsets. However, they typically require fine-tuning and their tracking performance for unknown time-varying trajectories is not guaranteed in general. Tracking of known reference trajectories with uncertain repetitive dynamics is studied in the context of iterative learning control (Bristow et al., 2006; Owens & Hätönen, 2005) with online extensions studying dynamic regret (Balta et al., 2022). However, the methods do not extend to non-repetitive setting of LQT with unknown time-varying references.
Predictive controllers such as model predictive control (MPC) have also been used to track constant references, see, for example, offset-free MPC (Pannocchia & Bemporad, 2007; Maeder et al., 2009). Various tracking MPC methods are proposed for time-varying references, e.g., (Köhler et al., 2020). However, unless we have an accurate prediction model for the reference trajectory, the guarantees do not extend to more general sequentially revealed unknown trajectories. Dual MPC methods with active exploration are proposed for tracking applications and uncertain dynamics (Heirung et al., 2017; Soloperto et al., 2019; Parsi et al., 2022). However, transient performance, regret guarantees, or the case of unknown time-varying references are not studied.
The LQT problem with unknown targets has been studied before from the point of view of control theory and adaptive control (Vamvoudakis, 2016; Modares & Lewis, 2014; Peterson & Narendra, 1982). Typically, the goal is to prove asymptotic convergence or boundedness of tracking error by appealing to Lyapunov stability theory. Most results assume time-invariant target dynamics. Tracking time-varying targets is possible using adaptive control techniques, under the assumption that the parameters of the target evolve like a random walk (Guo & Ljung, 1995). However, this assumption excludes adversarial targets. By using regret as a metric, we can obtain non-asymptotic guarantees that also capture adversarial, non-stochastic behaviors.
Finally, an alternative approach is to employ robust control techniques to account for worst-case disturbances/tracking errors. Notable approaches include the celebrated $\mathcal{H}_{\infty}$ control approach or mixed $\mathcal{H}_{2}-\mathcal{H}_{\infty}$ control (Zhang et al., 2021a). Recently, robust controllers inspired by the notion of regret were designed (Goel & Hassibi, 2023).
Appendix B Linear Quadratic Control Properties
@@ B.1 LQT properties
Let us now introduce the following proposition, which shows that the closed loop matrix $A-BK$ under the feedback gain $K$ defined in (8) is stable, that is, all of its eigenvalues are inside the unit circle. It also shows that the feedforward gains $K_{t},\forall t\leq T$ defined in (9) decay exponentially fast as $t$ increases. Under the optimal control law (7), we do not have stability in the classical sense; since the target state can be arbitrary the tracking error might not remain close to the origin. However, we have internal stability since all signals remain bounded as long as $r_{t}$ is bounded. A more general case with time-variant costs of the proposition is proven in Corollary 1 of (Zhang et al., 2021b).
Proposition B.1 (Stability (Zhang et al., 2021b)).
@@
$\|(A-BK)^{t}\|\leq\sqrt{\frac{\lambda_{max}(X)}{\lambda_{min}(X)}}\rho^{t},\,\rho=\sqrt{1-\frac{\lambda_{min}(Q)}{\lambda_{max}(X)}}\in(0,1)$
where $\lambda_{min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the minimum and maximum eigenvalue respectively, $X$ is defined in (6). As a result, the coefficient matrices $K_{t}$ defined in (9) satisfy
@@
for $c_{0}=\frac{\lVert B\rVert\lambda_{max}(X)}{\lambda_{min}(R)}\sqrt{\frac{\lambda_{max}(X)}{\lambda_{min}(X)}}$.
@@
Note that a pair $(Q,A)$ is detectable if and only if $(A^{\top},Q^{\top})$ is stabilizable. Under the above assumption, we still retain stability of $(A-BK)$ (Anderson & Moore, 2005). The only difference is that we do not have an explicit characterization of $\rho$ anymore as in Proposition B.1.
B.2 LQT as LQR with disturbances
We can recast the LQT problem into an LQR problem with adversarial disturbances, see, for example, (Karapetyan et al., 2023). By redefining the tracking error as the system state $e_{t}=x_{t}-r_{t}$ and encoding the time-varying target states as the disturbance $w_{t}=Ar_{t}-r_{t+1}$, the resulting LQR problem is given as,
$\min_{u_{0:T-1}}$ $\sum_{t=0}^{T-1}\left(\lVert e_{t}\rVert_{Q}^{2}+\lVert u_{t}\rVert_{R}^{2}\right)+\lVert e_{T}\rVert_{X}^{2}$ (15)
@@ B.2 LQT as LQR with disturbances
Appendix C Autoregressive Systems
@@
C.1 Representation and learning complexity
@@
Example C.1 (Constant Velocity Target).
@@
Example C.2 (Circular Target with Constant Speed).
@@
In many previous works (Karapetyan et al., 2023; Li et al., 2019; Nonhoff et al., 2023), learning complexity is captured by the total variation (path length) of the target itself
@@
In the case of ARX models we just replace $S_{t}$ with $\tilde{S}_{t}$. As shown in (Li et al., 2019, Th. 3), in the case of unstructured targets, e.g. randomly generated $r_{t}$, the former notion of complexity is optimal. However, we argue that in the case of dynamic structure, the former notion of complexity might be suboptimal. Consider the circular target or the linear target example. The total variation of the target state is linear with $T$ since $r_{t}$ is constantly changing: $L_{T}=\Omega(T)$. However, the total variation of the target dynamics is zero $V_{T}=0$. Using the PLOT algorithm would give us logarithmic regret in this case, while previous methods would give us linear regret.
We stress that every target can be represented trivially by ARX models. We can just set $S_{t}=0$, $v_{t}=r_{t}$. In this case, the total variation of $\tilde{S}_{t}=\begin{bmatrix}S_{t}&v_{t}\end{bmatrix}$ is equal to $L_{T}$, as in prior work. However, the representation is non-unique in general. In many cases, the underlying dynamic structure will imply that a lower complexity representation exists, e.g., see Examples C.1C.2. If the targets have no dynamic structure, then, our current setting still captures this case; the regret of PLOT will just be larger, i.e., of the same order as $L_{T}$. Hence, our dynamic regret bounds supersede the ones in prior work.
C.2 Multi-step ahead dynamics expressions